OpenStudy (faiithhh):

Which represents the solution to the absolute value equation? 5|2x - 4| + 1 = 11?

OpenStudy (faiithhh):

A) x = 1 or x = 3 B) x = 1 or x = -3 C) x = 3 or x = - 3 D) x = 6 or x = 1

OpenStudy (jiteshmeghwal9):

When 2x-4 >0 the equation becomes 5(2x-4)+1=11......(1) when 2x-4<0 the equation becomes -5(2x-4)+1=11 ........(2) solve equation (1)& (2) for x

OpenStudy (jiteshmeghwal9):

NOTE :- |x|=x for x>0 |x|=-x for x<0

OpenStudy (faiithhh):

am i just trying to find a replacement number for x?

OpenStudy (jiteshmeghwal9):

yes

OpenStudy (faiithhh):

okay, cause... 5(2x-4) + 1 = 11 2x - 4 = -2x 5(-2x)+ 1=11 -10x + 1 = 11 -9x = 11

OpenStudy (faiithhh):

right...?

OpenStudy (faiithhh):

so whatever x is has to get -9 up to 11?

OpenStudy (faiithhh):

im confused.

OpenStudy (jiteshmeghwal9):

5(2x+4)+1=11 subtract 1 from both sides 5(2x+4)=10 divide 5 from both sides (2x+4)=2 subtract 4 from both sides 2x=-2 divide by 2 both the side x=-1

OpenStudy (jiteshmeghwal9):

so u gt one solution for x as x=-1

OpenStudy (jiteshmeghwal9):

The other equation is -5(2x-4)+1=11 solve it similarly

OpenStudy (faiithhh):

OHHHH I SEE... -5(2x-4)+1=11 minus 1 both sides -5(2x-4)=10 divide by -5 both sides (2x-4) = -2 add 4 to both sides (2x) = 2 divide by two x = 1

OpenStudy (faiithhh):

i think...

OpenStudy (faiithhh):

@jiteshmeghwal9

563blackghost (563blackghost):

Actually I wouldnt agree....you first need to get the absolute equation to itself....first you must get `1` to the other side... \(\huge{5|2x-4|(1-1)=11-1}\) What would that equal?

OpenStudy (faiithhh):

5|2x-4|=10

563blackghost (563blackghost):

Correct :) Now we have `5|2x-4|=10`....now since 5 is `by` `|2x-4|` that would mean we would divide to get it to the other side...so we divide by 5... \(\huge{\frac{5}{5}|2x-4|=\frac{10}{5}}\) What would that equal?

OpenStudy (jiteshmeghwal9):

but i was correct though :P

OpenStudy (faiithhh):

|2x -4| = 2

563blackghost (563blackghost):

Not quite x does not equal to -1...

563blackghost (563blackghost):

Correct :) Now since this is a absolute equation the total can either be a positive or a negative sooo we would make two equation one with a positive total and one with a negative total.... \(\huge{2x-4=2}\) \(\huge{2x-4=-2}\) Do you understand?

OpenStudy (jiteshmeghwal9):

seriously there was a typo when i was solving my first equation i solved 5(2x+4)+1=11 instead of 5(2x-4)+1=11

OpenStudy (jiteshmeghwal9):

that's the reason behind why i gt x=-1 as my answer

563blackghost (563blackghost):

Oooo I see then yea technically you were correct :P It just i didnt recognize with the absolute and real numbers on both sides....I was wondering why you didnt get the absolute to itself...@jiteshmeghwal9

563blackghost (563blackghost):

@jiteshmeghwal9

OpenStudy (jiteshmeghwal9):

Haha nvm that was my mistake

563blackghost (563blackghost):

So we would simplify each equation to find x.... \(\huge{2x(-4+4)=(2+4)}\) We would subtract 4 from both sides....

563blackghost (563blackghost):

Sorry I meant we would add by 4 to cancel out `-4`...my mistake

563blackghost (563blackghost):

So we add it.... \(\huge{2x=6}\) Next to get the variable to itself we would divide by 2... \(\huge{\frac{2x}{2}=\frac{6}{2}}\) What would that equal?

OpenStudy (sshayer):

\[5\left| 2x-4 \right|+1=11,5\left| 2x+4 \right|=11-1=10\] \[\left| 2x-4 \right|=\frac{ 10 }{ 5 }=2\] \[2x-4=\pm 2\] \[2x-4=2,2x=2+4=6,x=\frac{ 6 }{ 2 }=3\] \[2x-4=-2,2x=-2+4=2,x=\frac{ 2 }{ 2 }=1\]